# Ordinality in Discrete-level Question Difficulty Estimation: Introducing Balanced DRPS and OrderedLogitNN

### v1

@@ 8 Conclusion
Experimental results indicate that OrderedLogitNN considerably outperforms existing methods on more complex tasks while performing comparably on simpler ones, both on probabilistic and degenerate predictions. Probabilistic predictions express uncertainty and are particularly valuable in downstream human-in-the-loop tasks such as selective prediction or active learning (Thuy and Benoit, 2023; Thuy et al., 2024). These findings underscore OrderedLogitNN as a highly attractive approach for discrete-level QDE. The learning dynamics of OrderedLogitNN deserve further investigation and future work can assess the impact of adapting the weight initializations and learning rate multiplier.

### v2

@@ 8 Conclusion
Experimental results indicate that OrderedLogitNN considerably outperforms existing methods on more complex tasks while performing comparably on simpler ones, both on probabilistic and degenerate predictions. Probabilistic predictions express uncertainty and are particularly valuable in downstream human-in-the-loop tasks such as selective prediction or active learning [29, 15]. These findings underscore OrderedLogitNN as a highly attractive approach for discrete-level QDE. The learning dynamics of OrderedLogitNN deserve further investigation and future work can assess the impact of adapting the weight initializations and learning rate multiplier.
@@ Acknowledgements.
Declaration on Generative AI
During the preparation of this work, the authors used GPT-4o in order to check grammar and spelling. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the publication’s content.